OpenStudy (anonymous):

Which expression is equivalent to two-thirds divided by five-sixths? five-sixths times two-thirds two-thirds times five-sixths three-halves times six-fifths two-thirds times six-fifths

OpenStudy (kc_kennylau):

two-thirds divided by five-sixths=(2/3)/(5/6)=(2/3)x(6/5)=D.